UT 20160423 LUCI1 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, APramskiy, DThompson,
Observer AO: DMiller, GTaylor (both LBTO), EPinna, APuglisi (remote from Italy)
Telescope Operator: SAllenson
LBTO Support: TSargent

Plan for the night

- Start with LUCI1-AO com, do variety of tests adapted to ambient conditions

- Bin3+4 tests by Doug & Greg

- Auto gain optimization LBT/Arcetri AO-teams

Summary

- AO-tests from LBTO/Arcetri-groups basically complete

- First encouraging AFC-tests on sky in K and Br_gam

- A few hickups at beginning but otherwise FLAO-system run reliably

- Mediocre to poor (1-2") seeing all night

Details

All times are UT
Files are luci1.20160423.NNNN.fits

02:00 Steve opens up. Weather looks good: Mostly clear, T = +9deg, Hum = 30%, Wind = 10 m/s

02:30 Steve checks pointing and collimates

PSF stability on short and intermediate timescales

02:40 OJ287a, preset OK, RunAO failed (deceleration error), resent preset, RunAO failed, restart AOS,
resent preset again, RunAO (bin=4, R=15mag) tooks too long (far more than 5min) LUCI receives and "error during commit"
TCS gave error since RunAO tooks so long (10min timeout), preset again, RunAO worked this time (Doug changed number
of iterations for bin=4, #2, DIMM=1.1-1.4", didn't execute the folloging offset, restartet scheduler, nope, restartet telescope service, nope
restart LUCI-SW, preset again, RunAO OK, scheduler problem, restart scheduler, finally started, #0003-0007, finally worked, IQ=0.22"
RunAO failed 3 times

04:01 PresetOJ287 _PSFa, OK, RunAO (bin=4, R=14.6), 04:07 taking data (6min overhead), #0008-0012,DIMM=1.2-1.4", IQ=0.15"
All worked fine

04:15 Preset OJ287b, OK, RunAO (bin=4), 04.23 data taking, #0013-0017, DIMM=1.2, IQ=0.14"
All worked fine

04:30 Preset OJ287 _PSFb, OK, RunA (bin=4), 04:37 data taking, #0018-0022, DIMM=1.1-1.2", IQ=0.15"
All worked fine

Duty Cycle 5/14mins = 36%, 5min on target = 5 dither * (20 * 3sec) integrated

Stability of system, low-surface brightness features

04:44 Preset HolmbergII, failed, corrected pointing, OK, RunAO (bin=2, R=12.7), 04:54 data taking
Br_gam: #0023, OK, loop opened (skip frame), #0024 crap, loop closed again, DIMM = 1.3-1.5", #0025 OK,
H2: #0026,0027 OK, FeII: #0028,0029 OK, N3.75: Br_gam #0030,0031 OK, H2: #0032,0033 OK , DIMM = 2" at the end
All worked fine Duty Cycle 41/63min = 65%

Seeing is pretty bad now. We are centered on a small high presure system and sandwiched by two low-pressure systems within 100km NW and SW frown, sad smile

Handover to Doug, Greg & Alfio, Enrico for optical gain tests

Log can be found at http://wiki.lbto.org/bin/view/AdaptiveOptics/20160423_AO_SX

Last 45min tests bin 3 and 4 by Doug & Greg

10:15 Cirrus moving in, DIMM constantly between 1 and 1.5"

Active flexure compensation

10:30 Preset to H1722_AFC, failed (entry for new N30 mask with 10 holes in database had an extra space), corrected, OK,
RunAO OK (bin=2, R=13.6), #0092-105 (K), #0106-0107 (Br_gam) took a bit to adjust everything, looks good!
First results on sky are encouraging

12:10 Steve's closing up

-- %USERSIG{JochenHeidt - 2016-04-23}%
Topic revision: r8 - 18 Oct 2016, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback